# Machine Learning-Based Estimation Of Wave Direction For Unmanned Surface Vehicles

### v1

@@ II Proposed model
Figure 1: Model overview (A temporal view)
@@ III-B Determining ground truth
In the controlled environment of the experimental pool, the wave direction is determined by the pool’s heading, since waves consistently originate from a fixed direction. For the open sea experiments, historical weather data are obtained from WindGuru [^1] https://www.windguru.cz/, which provides sea state information for the testing location.
@@ III-G Evaluation on unseen data
The proposed model is evaluated using sea data that were not included in the training phase to assess its generalization capabilities. A total of 136,041 raw data points are collected across 6 trajectories. Each trajectory lasts between 6 and 14 minutes. After applying the pre-processing steps described in Section II-A, the dataset is reduced to 134,350 data points.
The model’s angular error on these unseen data averages 40.34°. Figure 6 shows the predicted wave direction (after subtracting the drone’s yaw) for the first trajectory, where the ground truth dominant wave direction is 280.2°. When applying a moving average with a window of approximately 12 seconds, the predicted wave direction becomes notably more stable, as shown in Figure 7.
Figure 6: Predicted wave direction for open sea data (First trajectory)
Figure 7: Moving average on predicted wave direction for open sea data (First trajectory)
@@ V Conclusion
Future work will focus on expanding the dataset through further experiments in varied sea conditions and incorporating other maritime variables such as wind and current. This will help capture a broader range of scenarios and refine the model’s accuracy. Additionally, we plan to adapt the model for real-time applications, addressing challenges such as optimizing energy consumption and minimizing inference times. We also intend to combine our wave direction prediction model with a reinforcement learning agent to simulate and assess the USV’s behavior in complex and hostile environments, enhancing the robustness and autonomy of its navigation capabilities.

### v2

@@ II Proposed model
Figure 1: Model overview - A temporal view
@@ III-B Determining ground truth
In the controlled environment of the experimental pool, the wave direction is determined by the pool’s heading, since waves consistently originate from a fixed direction.
For the open sea experiments, the wave direction at the testing site is determined using traditional maritime estimation techniques (use of fixed landmarks).
@@ III-G Evaluation on unseen data
The proposed model is evaluated using sea data that were not included in the training phase to assess its generalization capabilities. A total of 136,041 raw data points are collected across six trajectories. Each trajectory lasts between 6 and 14 minutes. After applying the pre-processing steps described in Section II-A, the dataset is reduced to 134,350 data points.
The predicted dominant wave direction has a mean value of 314.64°. As noted in Section III-B, the wave direction was assessed using fixed landmarkers, resulting in an estimated direction of 335° with a tolerance of ± 5°. Based on this estimation, the model’s average error is determined to be 20.36°.
Figure 6 illustrates the predicted wave direction for the first trajectory (after subtracting the drone’s yaw). Applying a moving average with a window of approximately 16 seconds reduces the standard deviation to 13.78°, resulting in a notably more stable prediction, as depicted in Figure 7.
Figure 6: Predicted wave direction for open sea data - First trajectory
Figure 7: Moving average on predicted wave direction for open sea data - First trajectory
@@ V Conclusion
To further refine wave direction estimation, reliable field measurements of wave direction are needed (e.g., through the processing of images captured by an aerial drone). These measurements should be performed for each wave encountered by the drone, as the direction can fluctuate and evolve over time, in order to precisely measure the model’s error.
Future work will also focus on expanding the dataset through further experiments in varied sea conditions and incorporating other maritime variables such as wind and current. This will help capture a broader range of scenarios and refine the model’s accuracy. Additionally, we plan to adapt the model for real-time applications, addressing challenges such as optimizing energy consumption and minimizing inference times. We also intend to combine our wave direction prediction model with a reinforcement learning agent to simulate and assess the USV’s behavior in complex and hostile environments, enhancing the robustness and autonomy of its navigation capabilities.